OpenStudy (anonymous):

Pre-Calc: Need Help Fan + Medal for most helpful person! Eliminate the parameter. x = square root of t, y = 4t + 1 y = 4square root of x + 1, x ≥ 0 y = 4square root of x - 1, x ≥ 0 y = 4x2 - 1 y = 4x2 + 1

OpenStudy (anonymous):

\[x=\sqrt{t}\\x^2=t\] is a start

OpenStudy (anonymous):

so would it be the 4th answer? y= 4X^2 +1?

OpenStudy (loser66):

yup

OpenStudy (anonymous):

and does that make this one: Eliminate the parameter. x = t2 + 3, y = t2 - 2 y = x-5?

OpenStudy (loser66):

yup

OpenStudy (anonymous):

Do you know this last one? Eliminate the parameter. x = 5 cos t, y = 5 sin t

OpenStudy (loser66):

you get rid t of by square both sides of them, add them together and use trig cos^2 +sin^2 =1to simplify the stuff

OpenStudy (anonymous):

so how would we get both sin and cos in the same equation?

OpenStudy (loser66):

x^2 =25cos^2t y^2= 25sin^2t ---------------- x^2+y^2 =25 (cos^2 t + sin^2 t) =25 that is your final answer.

OpenStudy (anonymous):

thanks

OpenStudy (loser66):

ok